OpenStudy (anonymous):

5x-y=-6 -x+y=2 solve?

OpenStudy (anonymous):

is it moltiple chiece ?

myininaya (myininaya):

Add equations together first

OpenStudy (anonymous):

can you walk me through?

myininaya (myininaya):

(5x-y)+(-x+y)=(-6)+(2) Reorder terms 5x-x-y+y=-6+2 Combine like terms 4x+0=-4 4x=-4 so x equals what?

OpenStudy (anonymous):

...ihave no idea...

myininaya (myininaya):

How do you undo multiplication?

OpenStudy (anonymous):

divide

myininaya (myininaya):

so we need to divide by whatever is in front of x which is?

OpenStudy (anonymous):

4

myininaya (myininaya):

so if you divide both sides by 4 of 4x=-4 we get x=?

myininaya (myininaya):

-4/4?

myininaya (myininaya):

which is ?

myininaya (myininaya):

x=-4/4=-1

myininaya (myininaya):

now you can plug this into either equation you started with to find y